a has given loan to b of 52000 @15% rate of interest per annum for 3 years ,find the value of interest and the total amount b has to give back to a
1
Expert's answer
2019-11-25T12:32:22-0500

Its a simple interest question,

Thus Interest to be paid after 3 years,

"I=PRT=5200\times 0.15\times 3=2340"

Total Amount to be paid=Principal amount+ Interest=5200+2340=7540 unit